OpenStudy (hy123):

cot(theta)+1=0 what are all the solutions to theta?

OpenStudy (hy123):

So 180?

OpenStudy (anonymous):

Is it cos or cot??

OpenStudy (hy123):

cot

OpenStudy (anonymous):

And calc is using cos there..

OpenStudy (hy123):

Not sure.

OpenStudy (anonymous):

Oh sorry. I misread it.

OpenStudy (hy123):

Haha.

OpenStudy (anonymous):

\[\cot \theta = -1\] \[\frac{1}{\tan \theta} = -1\] \[\tan \theta = -1\] \[\tan \theta = \tan(\frac{- \pi}{4})\] General Solution of tanx = tana is given by: \[x = n \pi + a\] So, using this concept: \[\theta = n \pi + (\frac{- \pi}{4})\] \[\theta = n \pi - \frac{\pi}{4}\] \[\theta = \frac{(4n - 1)}{4}\pi\]

OpenStudy (hy123):

Thanks !

OpenStudy (anonymous):

Welcome dear..

OpenStudy (hy123):

Wait water. Is it possible if you could give me exact degrees to fit theta

OpenStudy (anonymous):

You can put n= 0, 1, 2, 3 there and can find theta according to that.. First one I do for you: Put n= 0, \[\theta = \frac{(4 \times 0 - 1)}{4}\pi = \frac{-\pi}{4}\] Like wise you can find more values of theta by putting n = 1, 2, 3 and so on..

OpenStudy (hy123):

Ahh okay thanks!